Is entropy a thermodynamic co ordinate? Please give me the answer. 

Grade:12th pass

2 Answers

Arun
5 years ago
The thermodynamic entropy therefore has the dimension of energy divided by temperature, and the unit joule per kelvin (J/K) in the International System of Units (SI). Thermodynamic entropy is an extensive property, meaning that it scales with the size or extent of a system.
Ad
5 years ago
Entropy H=nCp∆t
Internal energy U=nCv∆t
H=U+PV
Cp-Cv=R
Y=Cp/Cv
Q/t = KA(T1-T2)/x
According to Kirchoff's Law , the ratio of emissive power to that of absorptive power is the same for all bodies at a given temperature and it is equal to the emmisive power of a blackbody which is equal to 1.